Suppose that the firm output is given by q=aLK2 specify the marginal products of labour and capital


1
Expert's answer
2021-07-26T02:07:01-0400

For the firm output given,

the Marginal product of labour will be,

"\frac{\delta q}{\delta L}=1aL^{1-1}K^2 =\frac{q}{L}"

and the Marginal product of capital will be,

"\frac{\delta q}{\delta K}=2aLK^{2-1}=\frac{2q}{K}"